1. The Silver Zinc Cell represented as follows, can be used as a storage battery.
1. The Silver Zinc Cell represented as follows, can be used as a storage battery.
Zn(s)/ZnO(s)/KOH(aq., 40%), K2ZnO2(aq.)/Ag2O(s)/Ag(s)
a. Write the half-cell reactions, and the overall reaction
b. Calculate the standard cell potential of this cell.
c. If the actual cell potential is 1.70 V, what is the maximum possible energy density of this cell (kW-hr/Kg)?
1
Expert's answer
2020-05-14T14:34:14-0400

a. Write the half-cell reactions. At the negative plate (1.17 V):


"\text{Ag}_2\text{O} +2\text{H}^++2\text{e}^+\rightarrow2\text{Ag}+ \text{H}_2\text{O}."


In the electrolyte (-0.83 V):


"2\text{H}_2\text{O}\rightarrow2\text{H}^++2\text{OH}^-."

At the positive plate (1.22 V):


"\text{Zn}+2\text{OH}^-\rightarrow\text{Zn(OH)}_2+2\text{e}^-."


Write the overall reaction:


"\text{Zn}+\text{H}_2\text{O}+\text{Ag}_2\text{O}\rightarrow \text{Zn(OH)}_2+2\text{Ag}."

b. Calculate the standard cell potential of this cell:


"E=E_1+E_2+E_3=1.17-0.83+1.22=1.56\text{ V}."


c. The gravimetric capacity density for Zinc is 0.82 Ah/g. Therefore, the maximum possible energy density of this cell is:


"\omega=1.7\cdot0.82=1.394\text{ kWh/kg}."